You use 20.0 mL of solution A, 10.0 mL of solution B, and 80.0 mL of water for your first mixture. What is the initial concentration of KIO3? The stock solution A is 0.0200 M. Enter with the unit of M Your Answer:
Answer #1

initial concentration of KIO3 = 0.00364 M

Explanation

Total volume = 20.0 mL + 10.0 mL + 80.0 mL = 110.0 ml

According to dilution law, C1 * V1 = C2 * V2

C1 = concentration before mixing = 0.0200 M

V1 = volume before mixing = 20.0 mL

C2 = concentration after mixing

V2 = volume after mixing = 110.0 mL

C2 = C1 * (V1 / V2)

C2 = 0.0200 M * (20.0 mL / 110.0 mL)

C2 = 0.0200 M * (0.182)

C2 = 0.00364 M